OpenStudy (anonymous):

Suppose that 53% of us sleep between 7 and 8 hours a day. Express this percent as a decimal.

jimthompson5910 (jim_thompson5910):

x% = x/100 So for example, 23% = 23/100

jimthompson5910 (jim_thompson5910):

or 23% = 23/100 = 0.23 which means 23% = 0.23